OpenStudy (anonymous):

When finding solids of revolution (integration). Do you square y then integrate (then times by pi) or do you square the integral of pi (and then times by pi)?

OpenStudy (anonymous):

\[\pi \left( \int\limits_{a}^{b}y \right) ^{2}\] or...

OpenStudy (anonymous):

\[\pi \int\limits_{a}^{b}\left( y ^{2} \right)\]